OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. open parentheses x to the 3 fifths power close parentheses to the 2 thirds power

OpenStudy (dumbcow):

multiply the exponents

OpenStudy (anonymous):

\[(x^a)^b=x^{ab}\] \[(x^{\frac{3}{5}})^{\frac{2}{3}}=x^{\frac{3}{5}*\frac{2}{3}}\]